OpenStudy (anonymous):

help please? :o Ben is 8 years older than Daniel. Two years ago, Ben was 5 times older than Daniel.How old is Ben now?

OpenStudy (anonymous):

Dude where do you keep coming up with these crazy age questions?!

OpenStudy (anonymous):

its like my math hw ._.

OpenStudy (anonymous):

there's like a new one every second lol

OpenStudy (anonymous):

my teacher Dx

OpenStudy (anonymous):

weird math lol...

OpenStudy (anonymous):

lol yeah i have like 16 problems

OpenStudy (anonymous):

rightttt ._. my teacher doesnt help us at all

OpenStudy (anonymous):

Jeez lol and you can't figure out any ? :o

OpenStudy (anonymous):

no .-.

OpenStudy (anonymous):

they are kinda confusing lol..

OpenStudy (anonymous):

they are xD

OpenStudy (anonymous):

and annoying ._. but i mean i have to to them

OpenStudy (anonymous):

Your like "oh this looks easy" ... "NOPE"

OpenStudy (anonymous):

lmao yes!

OpenStudy (anonymous):

Hahaha Weird, I'd get my dad in here to figure these out but he's not home right now sorry x.x He's a pro at crazy questions lol

OpenStudy (anonymous):

lol "crazy questions"

OpenStudy (anonymous):

its okay tho xD

OpenStudy (anonymous):

You sir , seem to like my wording lol..

OpenStudy (anonymous):

im a girl xD lol and yes indeed

OpenStudy (anonymous):

You ma'am.*

OpenStudy (anonymous):

xD hah

OpenStudy (anonymous):

well best of luck to you!

OpenStudy (anonymous):

toodles!

OpenStudy (anonymous):

thank you c:

OpenStudy (anonymous):

Ben is 8 years older than Daniel. Two years ago, Ben was 5 times older than Daniel.How old is Ben now? B=D+8--(i) Two years ago, B-2 was Ben's age and D-2 Daniel's age Given--> B-2=5(D-2) --(ii) Now solve for B and D :) Answer should be Ben=12 years old now

OpenStudy (anonymous):

okay, thank you once again xD